Suppose a basket of goods and services has been selected to calculate the CPI and 2012 has been selected as the base year. In 2012, the basket’s cost was $77; in 2013, the basket’s cost was $82; and in 2014, the basket’s cost was $90. The value of the CPI in 2014 was

 
Answers: a.

116.9 and the inflation rate was 16.9%.

  b.

109.8 and the inflation rate was 9.8%.

  c.

109.8 and the inflation rate was 16.9%.
Answer #1

CPI 2014 = Price of basket in 2014/Price of basket in base year(2012)*100 = ($90/$77)*100 =116.9

CPI in 2013 = ($82/$77)*100 = 106.5

Inflation rate = (CPI in 2014 - CPI in 2013)/CPI in 2013*100 = [(116.9 - 106.5)/106.5]*100 = 9.8%